OpenStudy (anonymous):

Solve for x, showing the steps in your work. 5x + 9 -6 = 3x + 6 -3

OpenStudy (anonymous):

Can someone help me please?

OpenStudy (anonymous):

@Creepy

OpenStudy (anonymous):

I don't know how to solve this equation. :/

OpenStudy (anonymous):

what are your options?

OpenStudy (anonymous):

There are none.

OpenStudy (anonymous):

5x + 9 - 6 = 3x + 6 - 3 5x + 3 = 3x + 3 5x = 3x x=0

OpenStudy (anonymous):

You can give medals by clicking the 'Best Response' button.

OpenStudy (anonymous):

Okay. Um, why did you only subtract 3 from the right side and 6 from the left side?

OpenStudy (anonymous):

@Creepy

OpenStudy (anonymous):

9 - 6 = 3 6 - 3 = 3 5x + 3 = 3x + 3 5x = 3x

OpenStudy (anonymous):

Okay, thanks.